Two equal concurrent forces, each of F N have a resultant of 6 N. When the magnitude of one of the forces is doubled, the resultant becomes 11 N. Calculate the value of F and the angle between the forces.
1
Expert's answer
2020-06-25T18:33:47-0400

Let "F_1=F_2=x," "\ \ \alpha" - angle between these forces.

"\vec{F_3}=\vec{F_1}+\vec{F_2},\ \ F_3=6;"

"\vec{F_4}=2\vec{F_1}+\vec{F_2}=\vec{F_3}+\vec{F_1},\ \ F_4=11;"


We can find "F_3" using formula: "F_3=2F_2\cos(\alpha /2), \ \ 3=x\cos(\alpha /2)" .

("AB=BC, \ BD\perp AC\ \Rightarrow AC=AD= AB\cos(\alpha/2))"


And we can find "F_4" using law of cosines for "\Delta ECA:" "11^2=x^2+6^2-2\times x\times 6\times \cos (\pi-\alpha/2)=x^2+36+12x\cos(\alpha/2)=x^2+36+12\times 3,"

"121=72+x^2, \ \ x^2=49, \ \ x=7."

"3=x\cos(\alpha/2)=7\cos (\alpha/2), \ \ \cos(\alpha/2)=3/7, \ \ \alpha/2=\arccos 3/7, \ \ \alpha=2\arccos 3/7."


Answer: "F=7, \ \ \alpha =2\arccos 3/7."